# Bridging Earth and Space: A Survey on HAPS for Non-Terrestrial Networks

### v1

@@ III-E Energy and Sustainability
Energy efficiency is a critical consideration for future integrated TNs and NTNs. Traditional TNs rely on continuous operation of macro BSs to provide ubiquitous coverage, leading to substantial energy consumption even during low-traffic periods. This ”zero bit, non-zero watt” phenomenon reflects the inefficiency of maintaining coverage when demand is minimal.To address this issue, a promising solution involves offloading traffic from terrestrial macro BSs to NTN-powered hypercells, such as those supported by LEO satellites, HAPS, or tethered UAVs. During nighttime or periods of low demand, this approach allows power-intensive terrestrial macro BSs to enter low-power states or shut down entirely, significantly reducing energy consumption. In this context, LEO satellites are well-suited for offloading outdoor UEs due to their extensive coverage and ability to serve low-capacity traffic efficiently during nighttime. In contrast, HAPS, with their superior link budget and better signal penetration, are better suited for offloading indoor UEs, especially in urban environments where high spatial reuse is essential.
In [73], a heuristic algorithm was developed to optimize this offloading process. By prioritizing the least-utilized BS for offloading, and leveraging HAPS’s self-sufficient energy capabilities, the proposed approach minimized hourly energy consumption across the network. Simulation results indicated that this method could reduce energy consumption by up to 29% over a week, with savings reaching 41% during nighttime hours. These results highlight the complementary roles of HAPS in advancing sustainable and energy-efficient connectivity within integrated TNs and NTNs.
IV HAPS Architectures
@@
The integration of TNs and NTNs is critical to addressing the diverse connectivity challenges identified in Section III, such as extending coverage to remote areas, enabling robust disaster communication, managing unexpected traffic surges, supporting IoT applications, and facilitating advanced services such as mobile edge computing and AR/VR. By serving as dynamic platforms and connectors, HAPS can play a key role in bridging these systems and creating a resilient, high-capacity, global communication infrastructure.
@@
To this end, NTN platforms must efficiently transmit/receive communication signals between UEs, gateways, BS, and core networks. The NTN architecture, as defined by the 3GPP in [74], supports two fundamental payload designs: transparent (or bent-pipe) and regenerative. In the transparent approach, the NTN platform carries communication equipment that functions as a simple relay, forwarding signals without modification. In contrast, the regenerative architecture enables onboard signal processing at the NTN platform, allowing the digital processing of the —from packet data convergence protocol (PDCP) to physical (PHY) layer— to be performed directly within the platform. These architectural choices impact key factors such as spectral efficiency, rate, latency, and overall system complexity, shaping how NTNs integrate with TNs. Building on this foundation, we can distinguish three main network components within the NTN architecture:
  • Platforms, which carry communication equipment and implement either a transparent or regenerative payload to relay or generate communication signals.
  • Terminals, including direct UEs or specialized devices designed to interact with NTN platforms.
  • Gateways, which connect NTNs, e.g., satellites and HAPS, to core networks, facilitating data flow between TNs and NTNs.
At the same time, the NTN architecture incorporates the following types of links [75]:
  • Service links, which facilitate direct communication between NTN platforms and UEs, ensuring end-user connectivity across diverse environments.
  • Feeder links, which connect NTN platforms to gateways, operating at high frequencies to support high-capacity data transmission with error correction.
  • Backhaul links, which connect NTN gateways to core network facilities, typically relying on fiber-optic infrastructure to ensure high-speed data transport.
  • Inter-platform links, which enable communication between NTN platforms, supporting seamless handovers and reducing dependence on terrestrial infrastructure. They usually use FSO infrastructure.
@@
In Section IV-A, we will describe the different NTN components, including platforms, terminals, gateways, and their interactions, focusing on HAPS. In Section IV-B, we will discuss promising frequency bands for NTNs based on the link types, with considerations related to range, throughput, latency, and interference sensitivity. In Section IV-C, we will compare 3GPP bent-pipe and regenerative architectures, focusing on the role of HAPS. Finally, in Section IV-D, we will present relevant examples of field trials and deployments involving HAPS.
IV-A NTN Components
Table II: Comparison of HAPS platforms.
@@ IV-A NTN Components
In general, NTNs (see Fig. 1) include several key components, each playing a specific role in ensuring global, reliable communication. These components can be categorized into three main layers, namely space, aerial, and ground layers [79]. The space layer comprises geostationary equatorial orbit (GEO), MEO, and LEO satellites, while the aerial layer consists of HAPS and low-altitude UAVs. Satellites, HAPS, and UAVs in this case are considered as network platforms, which carry the node able to facilitate communications. However, in certain use cases, they may serve as network terminals, i.e., connectivity consumers, providing no communication functionality. The ground layer typically includes infrastructure that interacts with the space and aerial layers, such as gateway stations, core networks, and data centers. Considering TN-NTN integration, the different layers can be combined together to provide better coverage, capacity, resilience, and flexibility compared to standalone deployments [80]. Now, we shift our focus to the aerial layer, with a particular emphasis on HAPS, the subject of this survey. We distinguish between NTN platforms, terminals, and gateways.
IV-A1 Platforms
HAPS can be classified depending on the physical principle providing the lifting force:
  • Aerostatic platforms, which rely on buoyant force to stay aloft, and include airships and balloons.
  • Aerodynamic platforms, which generate lift through fixed or rotary wings to maintain flight.
  • Hybrid platforms, which rely on a combination of aerostatic and aerodynamic effects at different stages of flight, such as the tactical HAPS by CIRA [78].
All these types of HAPS have their own benefits and drawbacks, which we discuss in the following and are summarized in Table II.
@@ IV-A1 Platforms
Table III: Comparison of NTN terminals.
@@
Aerodynamic platforms offer significant advantages given their high flexibility, allowing them to move easily between regions in response to communication demands or environmental changes. This adaptability reduces the need for multiple airfields and enhances their capability for precise positioning through specific flight patterns. Modern aerodynamic designs are generally more robust against adverse weather conditions than aerostatic platforms, increasing their reliability in various operational scenarios. Additionally, many modern aerodynamic platforms, such as SoftBank’s “Sunglider” [77], can be entirely solar-powered, enabling them to remain operational for extended periods, ranging from weeks to months. However, these platforms come with some limitations. Their payload capacity is generally smaller compared to large aerostatic platforms, which can affect the scale and complexity of their communication systems. Moreover, the development and manufacturing costs of aerodynamic platforms are higher, reflecting their advanced engineering and specialized materials.
@@
IV-A2 Terminals
Platforms ultimately aim at delivering connectivity to network terminals, a.k.a, UEs. Given the broad spectrum of application scenarios outlined in Section III, it is evident that UEs extend far beyond conventional smartphones, and also include wearable devices, IoT sensors, robots, and UAVs. Terminals can be classified in terms of features like form factor, operational environment, and communication capabilities. In the following paragraphs, we will describe the main groups of terminals, namely portable consumer terminals, vehicular terminals, and fixed terminals, as well as their features, and specific requirements, as summarized in Table III.
@@ IV-A2 Terminals
Probably the largest group of UEs consists of portable consumer terminals, including smartphones, smartwatches, tablets, and laptops. These devices are designed for everyday use across urban, suburban, and rural areas, and require seamless handover between TNs and NTNs to ensure uninterrupted connectivity. The diverse range of capacity-demanding and latency-sensitive applications, such as gaming and video streaming, necessitates support for multiple frequency bands, compatibility with existing standards (e.g., fourth-generation (4G), 5G), and NTN-specific adaptations for satellite and HAPS links. Conventional mobile devices are further complemented by wearable and implantable terminals, such as fitness trackers, health monitors, and medical implants. These devices are typically ultra-portable, low-power, and designed for continuous operation. They often require only short-distance, low-data-rate communication with a UE’s hub device, such as a smartphone, relying on the ubiquitous coverage provided by the hub. Another significant group within portable terminals includes communication devices for first responders, disaster relief teams, and military personnel. These devices are specifically designed to operate in extreme conditions, providing reliable connectivity in areas where terrestrial infrastructure is damaged or nonexistent. They often require enhanced security and encryption features for sensitive communications, as well as high-capacity links to enable rapid data exchange in critical scenarios.
@@
Another distinct category of mobile devices comprises vehicular terminals, installed in cars, buses, trains, ships, and aircraft. These terminals operate under challenging conditions due to high mobility and diverse environments, including remote and maritime regions. Unlike handheld devices, they can accommodate more efficient power sources and advanced antennas, such as VSAT parabolic reflectors or phased array antennas.
very small aperture terminal (VSAT) systems typically feature parabolic dishes ranging from 0.75 m to 2.4 m in diameter, with some larger models, providing high-gain, narrow-beamwidth links ideal for point-to-point communications. The antenna size depends on required gain, operating frequency, and environmental factors. In contrast, phased arrays, such as those employed in Starlink or BGAN systems, are more compact and rely on electronic beamforming through controlled phase shifts across multiple radiating elements. This enables rapid, precise beam steering without mechanical movement, offering enhanced flexibility and robustness under vibration or harsh conditions. Modern designs integrate features like auto-acquisition, self-alignment, and automatic aiming, simplifying deployment and ensuring reliability during high-speed operations.
A related subset of vehicular devices includes unmanned vehicles, such as UAVs, self-driving cars, autonomous robots, and surveillance balloons. These systems combine high mobility with operational versatility across urban and remote areas. They require reliable, low-latency command-and-control links and high-capacity channels for real-time data transmission (e.g., video or telemetry). To maintain continuous connectivity beyond terrestrial coverage, NTN support—via platforms like LEO satellites or HAPS—is essential for ensuring global reach and service continuity.
@@
A distinctly different subset of fixed terminals comprises IoT devices, including sensors, indoor consumer smart household appliances and industrial equipment. While industrial equipment and household appliances can be connected through industrial gateways or indoor access points, remote sensors rely on their own connectivity solutions. These sensors are typically compact, low-power, and deployed in large numbers. They require massive machine type communication (mMTC) connectivity, often in remote or hard-to-reach locations, and energy-efficient communication protocols to maximize battery life.
This diverse range of UE requirements in integrated TN-NTN systems highlights the critical role of platform adaptability and technology integration. Meeting these diverse needs necessitates a collaborative ecosystem of terrestrial, satellite platforms, and HAPS, each uniquely suited to specific terminal categories and operating conditions. Specifically, the TN provides the best capacity and latency for densely populated areas, but is constrained in coverage and mobility support in remote regions, while satellites are indispensable for remote and hard-to-reach areas, but suffer from higher latency and lower rates. At the same time, HAPS serves as a versatile middle ground, offering good coverage, capacity, latency, and mobility, especially in semi-remote areas or regions requiring temporary connectivity.
IV-A3 Gateways
@@
IV-B NTN Links and Frequency Bands
Table IV: Comparison of frequency bands for NTN, in terms of range, throughput, latency, and interference sensitivity.
Band Range Throughput Latency Interference sensitivity
Sub-7 GHz 100s km Low (<<1 Gbps) Low Low
cm-Wave 10s–100 km Medium (1–10 Gbps) Medium Moderate (rain fade in Ku-band)
$m$-Wave 1–10 km High (10–100 Gbps) Medium High (severe rain fade)
Sub-THz <<5 km Ultra-high (100+ Gbps) Low Very high
THz <<1 km Tbps Very low Extreme
FSO 10s–100s km Ultra-high (Tbps) Very low Weather-dependent
@@ IV-B NTN Links and Frequency Bands
  • Service links provide connectivity between NTN platforms and UEs. These links must accommodate a diverse types of devices, ranging from stationary to highly mobile terminals, as described in Section IV-A, posing challenges such as signal power constraints, Doppler effects, and interference management.
@@
Notably, in Release 17 the 3GPP formally integrated NTN into the 5G new radio (NR) framework, defining use cases within the sub-7 GHz spectrum, particularly for massive access and narrowband IoT applications. Release 18 further expanded NTN support, introducing enhanced 5G NR operations for handheld terminals in the 10 GHz and above range. These updates focused on improved spectrum management and coexistence strategies to ensure seamless integration between TNs and NTNs [75, 74, 82].
The following subsections examine the key frequency bands currently used in NTN communications and those with potential for future adoption depending on the type of link, as summarized in Table IV.
IV-B1 Sub-7 GHz
@@
Referred to as 3GPP FR2, mmWave spans frequencies from 24.25 GHz to 71 GHz [83]. Given its large bandwidth and short orthogonal frequency division multiplexing (OFDM) symbols, this spectrum enables ultra-high data rates and extremely low latency, making it ideal for bandwidth-intensive applications such as enhanced mobile broadband (eMBB). mmWave also offers fiber-like speeds in fronthaul and backhaul communications, supporting network densification through small cells and high-directivity antennas.
Despite its advantages, mmWave has seen limited adoption in TNs due to its short range and high infrastructure costs. The severe free-space path loss, combined with high susceptibility to blockage from buildings, foliage, and human movement, requires a dense network of BSs, making large-scale deployment expensive. While small cell and advanced beamforming techniques based on MIMO helps to mitigate these limitations, deployment remains challenging, particularly in rural and remote areas [84].
@@
The use of mmWave in NTNs for broadband connectivity has recently gained significant research attention. The studies in [13, 85] explored NTN applications in 6G, evaluating aerial- and space-based mmWave links to provide connectivity for ground terminals under various network configurations. However, expanding NTN into mmWave bands requires effective spectrum-sharing techniques to balance isolation, licensing costs, and coexistence with existing TNs. Moreover, large-scale propagation models must be re-evaluated, considering factors such as path loss, penetration loss (also through the atmosphere), and shadow fading, which significantly impact NTN protocol design [86, 87].
IV-B3 cmWave
@@
Also known as 3GPP FR3, cmWave spans 7.125 GHz to 24 GHz, bridging the gap between sub-7 GHz and mmWave bands [88]. Unlike mmWave band, which suffer from severe propagation characteristics, FR3 offers improved coverage and efficiency, with channel properties closer to sub-7 GHz in the sub-12 GHz region than to mmWave [89, 90, 91, 92]. Propagation studies [93] indicate that FR3 communication at 10 GHz experiences approximately 9.5 dB higher path loss in LoS, and 9.7 dB in non-LoS, compared to 3.5 GHz; yet these losses remain manageable with advanced beamforming. This band also exhibits lower penetration loss than at mmWaves, making it a strong candidate for outdoor-to-indoor coverage [90]. Additionally, rain attenuation is negligible below 10 GHz, and reduced delay spread improves signal quality, while angular spreads remain consistent with sub-6 GHz deployments. A key challenge in FR3 networks is the coexistence with satellite services, particularly in the 12 GHz band, which is heavily utilized by LEO constellations such as Starlink [89]. Interference mitigation techniques, including advanced beamforming and ephemeris-based coordination, have been proposed to ensure harmonious spectrum sharing between terrestrial and satellite networks [90]. Dynamic spectrum access mechanisms and interference nulling strategies are also being explored to optimize spectral efficiency while maintaining reliable satellite communications [94].
IV-B4 Beyond-100 GHz: sub-THz and THz
The beyond-100 GHz region—covering sub-THz (92–174 GHz) and THz (0.1–10 THz)—offers extreme bandwidth for immersive communications and sensing [89, 95, 96]. Measurements at 140 GHz report 2.3 Gbps over 120 m, highlighting the rate–range trade-off and equalization needs [97]. Simulation studies also show that integrated sensing and communication (ISAC) in these bands can outperform 5G-spectrum baselines [98]. For NTN topologies, sub-THz has been proposed for backhaul between UAV-BSs and CPUs, with HAPS acting as the CPU to preserve LoS [99]. Feasibility analyses indicate that THz feeder/backhaul links around 10 Gbps are possible in light rainfall given careful link budgets and signal processing [100] Reconfigurable intelligent surfaces (reconfigurable intelligent surfaces) can further alleviate blockage [101].

### v2

@@ IV-E Energy and Sustainability
Energy efficiency is a critical consideration for future integrated TNs and NTNs. Traditional TNs rely on continuous operation of macro BSs to provide ubiquitous coverage, leading to substantial energy consumption even during low-traffic periods. This ”zero bit, non-zero watt” phenomenon reflects the inefficiency of maintaining coverage when demand is minimal. To address this issue, a promising solution involves offloading traffic from terrestrial macro BSs to NTN-powered hypercells, such as those supported by LEO satellites, HAPS, or tethered UAVs. During nighttime or periods of low demand, this approach allows power-intensive terrestrial macro BSs to enter low-power states or shut down entirely, significantly reducing energy consumption. In this context, LEO satellites are well-suited for offloading outdoor UEs due to their extensive coverage and ability to serve low-capacity traffic efficiently during nighttime. In contrast, HAPS, with their superior link budget and better signal penetration, are better suited for offloading indoor UEs, especially in urban environments where high spatial reuse is essential. Song et al. [201] developed a heuristic algorithm to optimize this offloading process. By prioritizing the least-utilized BS for offloading, and leveraging HAPS’s self-sufficient energy capabilities, the proposed approach minimized hourly energy consumption across the network. Simulation results indicated that this method could reduce energy consumption by up to 29% over a week, with savings reaching 41% during nighttime hours. These results highlight the complementary roles of HAPS in advancing sustainable and energy-efficient connectivity within integrated TNs and NTNs.
V HAPS Architectures
@@
The integration of TNs and NTNs is critical to addressing the diverse connectivity challenges identified in Section IV, such as extending coverage to remote areas, enabling robust disaster communication, managing unexpected traffic surges, supporting IoT applications, and facilitating advanced services such as mobile edge computing and AR/VR. To this end, NTN platforms must efficiently transmit/receive communication signals between UEs, gateways, BS, and core networks. While the recent progression of NTN standardization is focusing satellites [128], we expect, by serving as dynamic platforms and connectors, HAPS can play a key role in bridging NTN platforms and creating a resilient, high-capacity, global communication infrastructure.
V-A NTN Architecture
@@
NTNs refers to the networks which are using satellites or uncrewed aerial platforms in various constellations as relay nodes or BSs. In general, NTNs (see Fig. 2) include several key components, each playing a specific role in ensuring global, reliable communication. These components can be categorized into three main layers, namely space, aerial, and ground layers [96]. The space layer comprises LEO, MEO, GEO and even highly elliptical orbiting satellites, while the aerial layer consists of HAPS and low-altitude UAVs. Satellites, HAPS, and UAVs in this case are considered as network platforms, which carry the node able to facilitate communications. However, in certain use cases (for example, Earth observation satellites and HAPS or delivery drones), they may serve as network terminals, i.e., connectivity consumers, providing no communication functionality. The ground layer typically includes infrastructure that interacts with the space and aerial layers, such as gateway stations, core networks, and data centers. Considering TN-NTN integration, the different layers can be combined together to provide better coverage, capacity, resilience, and flexibility compared to standalone deployments [219]. The NTN architecture, as defined by the 3GPP in [2], supports two fundamental payload designs: transparent (or bent-pipe) and regenerative. In the transparent approach, the NTN platform carries communication equipment that functions as a simple relay, forwarding signals without modification. In contrast, the regenerative architecture enables onboard signal processing at the NTN platform, allowing the digital processing of the — from packet data convergence protocol (PDCP) to PHY layer — to be performed directly within the platform. These architectural choices impact key factors such as spectral efficiency, rate, latency, and overall system complexity, shaping how NTNs integrate with TNs. Building on this foundation, we can distinguish three main network components within the NTN architecture: (i) platforms, which carry communication equipment and implement either a transparent or regenerative payload to relay or generate communication signals; (ii) terminals, including direct UEs or specialized devices designed to interact with NTN platforms; (iii) gateways, which connect NTNs, e.g., satellites and HAPS, to core networks, facilitating data flow between TNs and NTNs.
@@
In Section V-B, we will describe the different NTN components, including platforms, terminals, gateways, and their interactions, focusing on HAPS. At the same time, the NTN architecture incorporates the following types of links [4]: service links, feeder links, backhaul links, and inter-platform links, which we discuss is subsection V-C, as well as the promising frequency bands for NTNs based on the link types, with considerations related to range, throughput, latency, and interference sensitivity. In Section V-D, we will compare 3GPP bent-pipe and regenerative architectures, focusing on the role of HAPS. Finally, in Section V-E, we will discuss ongoing and already closed HAPS-related projects, challenges they are facing and lessons which must be learned from the progress done so far.
V-B HAPS Platforms and Terminals
Table III: Comparison of HAPS platforms.
@@ V-B HAPS Platforms and Terminals
Now, we shift our focus to the aerial layer, with a particular emphasis on HAPS, the subject of this survey. We distinguish between NTN platforms, terminals, and gateways.
V-B1 Platforms
HAPS can be classified depending on the physical principle providing the lifting force: (i) aerostatic platforms, which rely on buoyant force to stay aloft, and include airships and balloons; (ii) aerodynamic platforms, which generate lift through fixed or rotary wings to maintain flight; (iii) hybrid platforms, which rely on a combination of aerostatic and aerodynamic effects at different stages of flight, such as the tactical HAPS by CIRA [104]. All these types of HAPS have their own benefits and drawbacks, which we discuss in the following and are summarized in Table III.
@@ V-B1 Platforms
Table IV: Comparison of NTN terminals.
@@
Aerodynamic platforms offer significant advantages given their high flexibility, allowing them to move easily between regions in response to communication demands or environmental changes. This adaptability reduces the need for multiple airfields and enhances their capability for precise positioning through specific flight patterns. Modern aerodynamic designs are generally more robust against adverse weather conditions than aerostatic platforms, increasing their reliability in various operational scenarios. Additionally, many modern aerodynamic platforms, such as SoftBank’s “Sunglider” [198], can be entirely solar-powered, enabling them to remain operational for extended periods, ranging from weeks to months. However, these platforms come with some limitations. Their payload capacity was generally smaller compared to large aerostatic platforms, which can affect the scale and complexity of their communication systems. Moreover, the development and manufacturing costs of aerodynamic platforms are higher, reflecting their advanced engineering and specialized materials.
@@
V-B2 Terminals
Platforms ultimately aim at delivering connectivity to network terminals, a.k.a, UEs. Given the broad spectrum of application scenarios outlined in Section IV, it is evident that UEs extends far beyond conventional smartphones, and also includes wearable devices, IoT sensors, robots, and UAVs. Terminals can be classified in terms of features like form factor, operational environment, and communication capabilities. In the following paragraphs, we will describe the main groups of terminals, namely portable consumer terminals, vehicular terminals, and fixed terminals, as well as their features, and specific requirements, as summarized in Table IV.
@@ V-B2 Terminals
Probably the largest group of UEs consists of portable consumer terminals, including smartphones, smartwatches, tablets, and laptops. These devices are designed for everyday use across urban, suburban, and rural areas, and require seamless handover between TNs and NTNs to ensure uninterrupted connectivity. The diverse range of capacity-demanding and latency-sensitive applications, such as gaming and video streaming, necessitates support for multiple frequency bands, compatibility with existing standards (e.g., 5G), and NTN-specific adaptations for satellite and HAPS links. Conventional mobile devices are further complemented by wearable and implantable terminals, such as fitness trackers, health monitors, and medical implants. These devices are typically ultra-portable, low-power, and designed for continuous operation. They often require only short-distance, low-data-rate communication with a UE’s hub device, such as a smartphone, relying on the ubiquitous coverage provided by the hub. Another significant group within portable terminals includes communication devices for first responders, disaster relief teams, and military personnel. These devices are specifically designed to operate in extreme conditions, providing reliable connectivity in areas where terrestrial infrastructure is damaged or nonexistent. They often require enhanced security and encryption features for sensitive communications, as well as high-capacity links to enable rapid data exchange in critical scenarios.
@@
Another distinct category of mobile devices comprises vehicular terminals, installed in cars, buses, trains, ships, and aircraft. These terminals operate under challenging conditions due to high mobility and diverse environments, including remote and maritime regions. Unlike handheld devices, they can accommodate more efficient power sources and advanced antennas, such as VSAT parabolic reflectors or phased array antennas. Typically, very small aperture terminal (VSAT) systems feature parabolic dishes ranging from 0.75 m to 2.4 m in diameter, with some larger models, providing high-gain, narrow-beamwidth links ideal for point-to-point communications. The antenna size depends on required gain, operating frequency, and environmental factors. In contrast, phased arrays, such as those employed in Starlink or BGAN systems, are more compact and rely on electronic beamforming through controlled phase shifts across multiple radiating elements. This enables rapid, precise beam steering without mechanical movement, offering enhanced flexibility and robustness under vibration or harsh conditions. Modern designs integrate features like auto-acquisition, self-alignment, and automatic aiming, simplifying deployment and ensuring reliability during high-speed operations. A related subset of vehicular devices includes unmanned vehicles, such as UAVs, self-driving cars, autonomous robots, and surveillance balloons. These systems combine high mobility with operational versatility across urban and remote areas. They require reliable, low-latency command-and-control links and high-capacity channels for real-time data transmission (e.g., video or telemetry). To maintain continuous connectivity beyond terrestrial coverage, NTN support—via platforms like LEO satellites or HAPS—is essential for ensuring global reach and service continuity.
@@
A distinctly different subset of fixed terminals comprises IoT devices, including sensors, indoor consumer smart household appliances and industrial equipment. While industrial equipment and household appliances can be connected through industrial gateways or indoor access points, remote sensors rely on their own connectivity solutions. These sensors are typically compact, low-power, and deployed in large numbers. They require massive machine type communication (mMTC) connectivity, often in remote or hard-to-reach locations, and energy-efficient communication protocols to maximize battery life. This diverse range of UE requirements in integrated TN-NTN systems highlights the critical role of platform adaptability and technology integration. Meeting these diverse needs necessitates a collaborative ecosystem of terrestrial, satellite platforms, and HAPS, each uniquely suited to specific terminal categories and operating conditions. Specifically, the TN provides the best capacity and latency for densely populated areas, but is constrained in coverage and mobility support in remote regions, while satellites are indispensable for remote and hard-to-reach areas, but suffer from higher latency and lower rates. At the same time, HAPS serves as a versatile middle ground, offering good coverage, capacity, latency, and mobility, especially in semi-remote areas or regions requiring temporary connectivity.
V-B3 Gateways
@@
V-C NTN Links and Frequency Bands
Table V: Comparison of frequency bands for NTN, in terms of range, throughput, latency, and interference sensitivity.
Band Typical range Peak throughput Latency profile Interference sensitivity
Sub-7 GHz (410 - 7125 MHz) Hundreds of km Low (<<1 Gbps) Low Low (robust propagation)
cm-Wave (7.125 - 24.25 GHz) Tens–hundreds of km Medium (1–10 Gbps) Medium Moderate (rain fade in Ku-band)
mm-Wave (24.25 - 52.6 GHz), FR2-2 (52,6 - 71 GHz) <<10 km High (10–100 Gbps) Medium High (severe rain attenuation)
Sub-THz (100 - 300 GHz) <<5 km Ultra-high (100+ Gbps) Low Very high (strong atmospheric loss)
THz (0.1 - 10 THz) <<1 km Tbps-class Very low Extreme (highly fragile links)
FSO (3 THz - 30 PHz) Up to hundreds of km Ultra-high (Tbps) Very low Weather-dependent (fog/clouds)
@@ V-C NTN Links and Frequency Bands
  • Service links provide connectivity between NTN platforms and UEs. These links must accommodate a diverse types of devices, ranging from stationary to highly mobile terminals, as described in Section V-B, posing challenges such as signal power constraints, Doppler effects, and interference management.
@@
Notably, in Release 17 the 3GPP formally integrated NTN into the 5G NR framework, defining use cases within the sub-7 GHz spectrum, particularly for massive access and narrowband IoT applications. Release 18 further expanded NTN support, introducing enhanced 5G NR operations for handheld terminals in the 10 GHz and above range. These updates focused on improved spectrum management and coexistence strategies to ensure seamless integration between TNs and NTNs [4, 2, 1].
The following subsections examine the key frequency bands currently used in NTN communications and those with potential for future adoption depending on the type of link, as summarized in Table V.
V-C1 Sub-7 GHz
@@
Referred to 3GPP FR2-1, mmWave spans frequencies from 24.25 GHz to 52.6 GHz while 3GPP FR2-2 designated for frequencies from 52.6 GHz to 71 GHz [8].Given its large bandwidth and short orthogonal frequency division multiplexing (OFDM) symbols, this spectrum enables ultra-high data rates and extremely low latency, making it ideal for bandwidth-intensive applications such as enhanced mobile broadband (eMBB). mmWave also offers fiber-like speeds in fronthaul and backhaul communications, supporting network densification through small cells and high-directivity antennas.
@@
Despite its advantages, mmWave has seen limited adoption in TNs due to its short range and high infrastructure costs. The severe free-space path loss, combined with high susceptibility to blockage from buildings, foliage, and human movement, requires a dense network of BSs, making large-scale deployment expensive. While small cell and advanced beamforming techniques based on MIMO helps to mitigate these limitations, deployment remains challenging, particularly in rural and remote areas [45]. The use of mmWave in NTNs for broadband connectivity has recently gained significant research attention. The studies [77, 78] explored NTN applications in 6G, evaluating aerial- and space-based mmWave links to provide connectivity for ground terminals under various network configurations. However, expanding NTN into mmWave bands requires effective spectrum-sharing techniques to balance isolation, licensing costs, and coexistence with existing TNs. Moreover, large-scale propagation models must be re-evaluated, considering factors such as path loss, penetration loss (also through the atmosphere), and shadow fading, which significantly impact NTN protocol design [178, 190].
V-C3 cmWave
@@
Also known as 3GPP FR3, cmWave spans 7.125 GHz to 24.25 GHz, bridging the gap between sub-7 GHz and mmWave bands [5]. Unlike mmWave band, which suffer from severe propagation characteristics, FR3 offers improved coverage and efficiency, with channel properties closer to sub-7 GHz in the sub-12 GHz region than to mmWave [52, 115, 154, 193]. Propagation studies [99] indicate that FR3 communication at 10 GHz experiences approximately 9.5 dB higher path loss in LoS, and 9.7 dB in non-LoS, compared to 3.5 GHz; yet these losses remain manageable with advanced beamforming. This band also exhibits lower penetration loss than at mmWaves, making it a strong candidate for outdoor-to-indoor coverage [115]. Additionally, rain attenuation is negligible below 10 GHz, and reduced delay spread improves signal quality, while angular spreads remain consistent with sub-6 GHz deployments. A key challenge in FR3 networks is the coexistence with satellite services, particularly in the 12 GHz band, which is heavily utilized by LEO constellations such as Starlink [52]. Interference mitigation techniques, including advanced beamforming and ephemeris-based coordination, have been proposed to ensure harmonious spectrum sharing between terrestrial and satellite networks [115]. Dynamic spectrum access mechanisms and interference nulling strategies are also being explored to optimize spectral efficiency while maintaining reliable satellite communications [114].
V-C4 Beyond-100 GHz: sub-THz and THz